# Weak type (1,1) jump inequalities in a nonsymmetric Gaussian setting

### v1

@@ 1. Introduction
If $\mathcal{I}\subset\mathbb{R}_{+}$ is an interval and $f\in L^{1}(\gamma_{\infty})$, the $\varrho$-th order variational seminorm of $\mathcal{H}_{t}f(x)$ on $\mathcal{I}$ is defined by
@@
If $\mathcal{I}\subset\mathcal{R}_{+}$ is an interval, $\lambda>0$ and $f\in L^{1}(\gamma_{\infty})$, the $\lambda$-jump counting function for $\mathcal{H}_{t}\,f$ in $\mathcal{I}$ at $x\in\mathbb{R}^{n}$ is defined as
@@
$J_{2}^{1,\infty}(\mathcal{H}_{t}f:\,t>0)\lesssim\|f\|_{L^{1}(\gamma_{\infty})}.$ (1.8)
@@
We will write $C<\infty$ and $c>0$ for various constants, all of which depend only on $n$, $Q$ and $B$, unless otherwise explicitly stated. The Lebesgue measure of a measurable set $E\subset\mathbb{R}^{n}$ will be denoted by $|E|$. We write $B(x_{0},r)$ for the closed ball of center $x_{0}$ and radius $r$, and $CB(x_{0},r)$ for its concentric scaling. By $\mathbb{N}$ we mean $\{0,1,\dots\}$.
@@ 2. Subadditivity
Let $\varphi=\varphi(t)$ be a continuous real- or complex-valued function defined in an interval $\mathcal{I}\subset\mathbb{R}_{+}$. The $\lambda$-jump counting function of $\varphi$ in $\mathcal{I}$ may be defined as in (1.2), with $\varphi(t)$ replacing $\mathcal{H}_{t}f(x)$. Then $N_{\lambda}\big(\varphi(t):\,t\in\mathcal{I}\big)$ is subadditive in $\mathcal{I}$. Moreover, it satisfies a kind of quasi-subadditivity property with respect to $\varphi$; more specifically, one has
@@ 5. Localization
$J^{1,\infty}_{\varrho}\left(\mathcal{H}_{t}f:t\in[1,\infty)\right)\lesssim\|f\|_{L^{1}(\gamma_{\infty})}$ (5.6)
@@
$J^{1,\infty}_{\varrho}\left(\mathcal{H}_{t}^{\mathrm{glob}}f:t\in(0,1]\right)\lesssim\|f\|_{L^{1}(\gamma_{\infty})}.$ (5.7)
@@ 6. The local case for $t$ not too small
We can safely apply (4.6) to the operator $f\mapsto\widetilde{r}_{j}\,\mathcal{H}_{t}\,(f\,r_{j})$. This was verified for $\mathcal{H}_{t}$ in [13, Subsection 5.1], and the extra factors $\widetilde{r}_{j}$ and $r_{j}$ do not matter here. Combined with (6.2) and (6.1), this yields for any $\rho\geq 1$ and $f\in L^{1}(\gamma_{\infty})$
@@ 7. The local case for small $t$ : the decomposition
In this and the following sections, $j\in\mathbb{N}$ will be fixed. We consider the case when $0<t\leq\min(1,|x_{j}|^{-2})$. Then $t\lesssim 1/|x|^{2}$, as seen by considering small and large $x_{j}$ separately. The operator $\mathcal{H}_{t}^{\mathrm{loc}}$ will be decomposed into a sum of three difference operators and a main operator. More precisely, we write
@@
The critical issues are concentrated in $M_{t}$, and one notices that this operator basically coincides, up to some constants and the factor $e^{R(x)}$, with the operator with kernel $K_{t}^{c}(x-u)$ introduced in [15, formula (8.3)]. The results in [15, Section 8] imply that the variation operator
@@
is not of strong nor weak type $(p,p)$ with respect to $\gamma_{\infty}$, for any $p\in[1,\infty)$. However, we shall prove in Section 9 that jump quasi-seminorms with $\varrho=2$ behave well for $M_{t}$.
Before considering $M_{t}$, we study the three difference operators in the next section, proving that they satisfy good variational inequalities for all $\varrho\geq 1$.
@@ 8.1. Some auxiliary bounds
$\lesssim\frac{e^{R(x)}}{t^{n/2}}\,e^{R(x)}\exp\left(-c\,\frac{|u-x|^{2}}{t}\right)\,\left(\frac{|u-x|^{2}}{t}+\frac{|u-x|^{4}}{t^{2}}\right)\,$
@@ 8.2. Boundedness of the variation operator for the differences
$C+\log_{+}\frac{\min(1,|x_{j}|^{-2})}{|u-x|^{2}}.$
We insert these estimates in the last double integral in (8.28), after observing that there $u-x\in{4B_{j}}-{6B_{j}}\subset B(0,10/(1+|x_{j}|))$ and also $e^{R(u)}\simeq e^{R(x)}\simeq e^{R(x_{j})}$. Thus we can replace $d\gamma_{\infty}(u)$ by $du$ and delete the factor $e^{R(x)}$, in this integral. Further, $1+|x|\lesssim 1+|x_{j}|$ there. For $n>1$ we get
@@ 9. The local case for small $t$ : the main operator $M_{t}$
$J_{2}^{p}\left(\widetilde{r}_{j}\,M_{t}(f\,r_{j}):0<t\leq\min(1,|x_{j}|^{-2})\right)\lesssim\|f\,r_{j}\|_{L^{p}(\gamma_{\infty})}$ (9.1)
@@
$\simeq e^{R(x_{j})}\,e^{-R(x_{j})/p}\left\lVert\,ce^{-R(x_{j})}\lambda\sqrt{N_{c\,\lambda e^{-R(x_{j})}}\left(\big(\psi_{t}*\big(fr_{j}e^{-R(\cdot)}\big)\right)}\,\right\rVert_{L^{p}(6B_{j},dx)}$

### v2

@@ 1. Introduction
[…]
@@
[…]
@@
[…]
with $C=C(n,Q,B)$.
@@
We will write $C<\infty$ and $c>0$ for various constants, all of which depend only on $n$, $Q$ and $B$, unless otherwise explicitly stated. The constants involved when we write $\lesssim$ or $\gtrsim$ or $\simeq$ are of this type, except when there is a subscript indicating dependence also on an additional parameter. The Lebesgue measure of a measurable set $E\subset\mathbb{R}^{n}$ will be denoted by $|E|$. We write $B(x_{0},r)$ for the closed ball of center $x_{0}$ and radius $r$, and $CB(x_{0},r)$ for its concentric scaling. By $\mathbb{N}$ we mean $\{0,1,\dots\}$.
@@ 2. Subadditivity
[…]
@@ 5. Localization
$J^{1,\infty}_{\varrho}\left(\mathcal{H}_{t}f:t\in[1,\infty)\right)\lesssim_{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\varrho}\|f\|_{L^{1}(\gamma_{\infty})}$ (5.6)
@@
$J^{1,\infty}_{\varrho}\left(\mathcal{H}_{t}^{\mathrm{glob}}f:t\in(0,1]\right)\lesssim_{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\varrho}\|f\|_{L^{1}(\gamma_{\infty})}.$ (5.7)
@@ 6. The local case for $t$ not too small
We can safely apply (4.6) to the operator $f\mapsto\widetilde{r}_{j}\,\mathcal{H}_{t}\,(f\,r_{j})$. This was verified for $\mathcal{H}_{t}$ in [13, Subsection 5.1], and the extra factors $\widetilde{r}_{j}$ and $r_{j}$ do not matter here. Combined with (6.2) and (6.1), this yields for any $\varrho\geq 1$ and $f\in L^{1}(\gamma_{\infty})$
@@ 7. The local case for small $t$ : the decomposition
In this and the following sections, we keep $j\in\mathbb{N}$ fixed and prove Proposition 5.2 in the remaining case $0<t\leq\min(1,|x_{j}|^{-2})$. For these $t$ one has $t\lesssim 1/|x|^{2}$, as seen by considering small and large $x_{j}$ separately.
The operator $\mathcal{H}_{t}^{\mathrm{loc}}$ will be decomposed into a sum of three difference operators and a main operator. More precisely, we write
@@
In the next section, we shall see that the three difference operators satisfy good variational inequalities for all $\varrho\geq 1$. Thus the critical part of $\mathcal{H}_{t}^{\mathrm{loc}}$ or $\mathcal{H}_{t}$ is $M_{t}$. One notices that $M_{t}$ basically coincides, up to some constants and the factor $e^{R(x)}$, with the operator with kernel $K_{t}^{c}(x-u)$ introduced in [15, formula (8.3)]. The results in [15, Section 8] therefore imply that the variation operator
@@
is not of strong nor weak type $(p,p)$ with respect to $\gamma_{\infty}$, for any $p\in[1,\infty)$. It follows that the same negative result holds for $\mathcal{H}_{t}^{\mathrm{loc}}$ and $\mathcal{H}_{t}$.
However, we shall prove in Section 9 that jump quasi-seminorms with $\varrho=2$ behave well for $M_{t}$, in the way stated in Theorem 1.2.
Before considering $M_{t}$, we study the three difference operators.
@@ 8.1. Some auxiliary bounds
$\lesssim\frac{e^{R(x)}}{t^{n/2}}\,{\cancel{{e^{R(x)}}}}\exp\left(-c\,\frac{|u-x|^{2}}{t}\right)\,\left(\frac{|u-x|^{2}}{t}+\frac{|u-x|^{4}}{t^{2}}\right)\,$
@@ 8.2. Boundedness of the variation operator for the differences
$C+\log_{+}\frac{\min(1,|x_{j}|^{-2})}{|u-x|^{2}},$
where $\log_{+}\beta=\max(0,\beta)$ for $\beta>0$. We insert these estimates in the last double integral in (8.28), after observing that there $u-x\in{4B_{j}}-{6B_{j}}\subset B(0,10/(1+|x_{j}|))$ and also $e^{R(u)}\simeq e^{R(x)}\simeq e^{R(x_{j})}$. Thus we can replace $d\gamma_{\infty}(u)$ by $du$ and delete the factor $e^{R(x)}$, in this integral. Further, $1+|x|\lesssim 1+|x_{j}|$ there. For $n>1$ we get
@@ 9. The local case for small $t$ : the main operator $M_{t}$
$J_{2}^{p}\left(\widetilde{r}_{j}\,M_{t}(f\,r_{j}):0<t\leq\min(1,|x_{j}|^{-2})\right)\lesssim_{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}p}\|f\,r_{j}\|_{L^{p}(\gamma_{\infty})}$ (9.1)
@@
$\simeq_{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}p}e^{R(x_{j})}\,e^{-R(x_{j})/p}\left\lVert\,ce^{-R(x_{j})}\lambda\sqrt{N_{c\,\lambda e^{-R(x_{j})}}\left(\big(\psi_{t}*\big(fr_{j}e^{-R(\cdot)}\big)\right)}\,\right\rVert_{L^{p}(6B_{j},dx)}$